OpenStudy (anonymous):

if PR is 4x-10 and RQ is 2x+12 whats PR?

OpenStudy (anonymous):

okay do you have like a picture:)

OpenStudy (anonymous):

and are they congruent

OpenStudy (anonymous):



OpenStudy (anonymous):

yes they are congruent! okay so all you have to do is solve for x..4x-10=2x+12

OpenStudy (anonymous):

yh and x would be 11, but what would be PR

OpenStudy (anonymous):

plug in the x into your answer

OpenStudy (anonymous):

so would it be right if i said that 4x11-10? that would equal 44-10= 33 (PR=33)